# Is the age pension in Australia sustainable and fair? Evidence from forecasting the old-age dependency ratio using the Hamilton-Perry model

### v1

@@ Abstract
The age pension aims to assist eligible elderly Australians meet specific age and residency criteria in maintaining basic living standards. In designing efficient pension systems, government policymakers seek to satisfy the expectations of the overall aging population in Australia. However, the population’s unique demographic characteristics at the state and territory level are often overlooked due to the lack of available data. We use the Hamilton-Perry model, which requires minimum input, to model and forecast the evolution of age-specific populations at the state level. We also integrate the obtained sub-national demographic information to determine sustainable pension ages up to 2051. We also investigate pension welfare distribution in all states and territories to identify disadvantaged residents under the current pension system. Using the sub-national mortality data for Australia from 1971 to 2021 obtained from AHMD (2023), we implement the Hamilton-Perry model with the help of functional time series forecasting techniques. With forecasts of age-specific population sizes for each state and territory, we compute the old age dependency ratio to determine the nationwide sustainable pension age.
@@ 1 Introduction
We propose modeling and projecting state and territory populations using the Hamilton-Perry (HP) method (Hamilton & Perry, 1962) and use the forecasts to determine the sustainable pension age. The HP method overcomes the restriction of data availability at the sub-national level since it requires significantly fewer inputs and computational time to implement than the traditional cohort-component method (see, e.g., Shang et al., 2016) or multiregional population projection (see, e.g., Rogers, 1975). Specifically, the standard cohort-component approach for population projection requires high-quality data about mortality, fertility, and immigration of a population and assumptions. In contrast, the HP model only requires population size by age and sex from the two most recent censuses (Smith et al., 2001, Chapter 7). This has made the HP method popular over the last decade in small area population forecasts, with notable applications including an initial population projection for Clark County, Nevada (Swanson et al., 2010) and a subsequent nationwide census tract evaluation across all sub-county areas in the United States (Baker et al., 2021), to name only a few. Most recently, a modified HP method was considered to forecast populations for small areas of Australia at the statistical area Level 2 scale (defined as in Australian Bureau of Statistics, 2011). This modified HP method with smoothed age profiles produces the most accurate forecasts (Wilson & Grossman, 2022).
This paper contributes a new variant of the HP method by extending the idea of a smoothing cohort change ratio (CCR, Wilson & Grossman, 2022) in combination with functional time series forecasting techniques (see Ramsay & Silverman, 2005; Ferraty & Vieu, 2006, for an introduction to functional data analysis). Functional time series methods have been widely used in demographic data modeling and forecasting (see, e.g., Cairns et al., 2011; D’Amato et al., 2011; Carfora et al., 2017). Using this framework, the CCR of sex $s$ increasing from age $x$ to $x+1$ within the time interval $[t-1,t]$, $\text{CCR}^{s}_{x\rightarrow x+1}(t-1\rightarrow t)$, can be regarded as a stochastic process defined on a compact set $x\in\mathcal{I}$ and indexed by a time label $t\in\mathbf{Z}$. This functional time series setting is best known for retaining the correlation of demographic observations at different ages, which enables extracting the temporal dependence of the underlying stochastic process for forecasting purposes. These advantages allow functional time series methods to produce more accurate forecasts than the commonly used parametric models (see, e.g., Hyndman & Ullah, 2007). With minimal data input, we combine functional time series forecasting techniques with the HP model to produce accurate sub-national population forecasts. By aggregating the sub-national population forecasts to the national level, we determine a sustainable pension age based on the current old-age dependency ratio.
@@
The remainder of the paper is organized as follows. Section 2 describes the age- and sex-specific population data at Australia’s national and sub-national levels. Section 3 provides details of the population forecasting method. The empirical data analysis results are presented in Section 4. Section 5 proposes an approach to determine the optimal pension age based on the forecasted population and its old-age dependency ratio. With the help of the HP model, we can study the changes in population at the sub-national level. Based on the historical observed data, we generate out-of-sample population forecasts from 2022 to 2051 and subsequently determine a sustainable pension age for Australia. With the sub-national population projections and forecasts of pension payout rates, we compare lifetime age pension income for residents in eight states and territories in Australia in Section 6. Finally, Section 7 concludes the paper and provides some directions for future research.
@@ 2.2 Sub-national age- and sex-specific data
While national-level data provides an overall picture of a country’s demographic trends, economic indicators, and other key statistics, sub-national details are crucial for understanding regional variations and making more targeted policy decisions. The sub-national data allow for granular analysis of regional disparities, urban-rural differences, resource allocation, and infrastructure planning.
@@ 3.1 Hamilton-Perry model
A recent study by Wilson & Rees (2021) outlined the standard process for creating a national population forecast using a cohort-component model. This model requires high-quality data on fertility, mortality, and internal and international migration. The quality of these datasets needs to be carefully examined to ensure accurate population forecasts. The cohort-component model is advantageous as it effectively represents underlying population processes and provides age-disaggregated projections for various purposes such as government planning, policy making, budgeting, and service delivery. However, it can also be highly inaccurate if poor assumptions are made about fertility, mortality, and migration. Additionally, it requires significant input data, and assumptions about future fertility, mortality, and migration must be prepared separately.
Compared to the cohort-component model, Hamilton & Perry’s (1962) method overcomes these drawbacks as it requires much smaller data requirements for projection. Instead of modeling mortality, fertility, and migration, the HP method requires data from the two most recent censuses (Smith et al., 2001, pp. 153–158). The ratio between two population censuses is known as the CCR. It measures the combined net effect of mortality and migration experienced by the cohort over time. It is estimated from population sizes by age and sex for the jump-off year and (commonly) one year earlier. For example, the CCR for the cohort of sex $s$ at age interval $x$ at time $t-1$ and age $x+1$ at time $t$ is defined as
@@
The HP method is commonly used for predicting the population of small areas (e.g., Wilson et al., 2022). With the sub-national population forecasts for eight regions, we study the dynamic changes in regional population counts. By aggregating the population forecasts at the national level, we provide Australian Government policymakers with a future sustainable retirement age.
@@ 3.2 Functional time series forecasting
We treat a time series of age-specific populations as a functional time series. Define $\mathcal{X}$ as a stochastic process defined on a compact set $\mathcal{I}$, with finite variance $\int_{\mathcal{I}}E(\mathcal{X}^{2})<\infty$. Let $\mu$ be the mean function of $\mathcal{X}$, the covariance function of $\mathcal{X}$ is defined as
@@
There are many ways of determining the optimal number of $K$, such as the bootstrap approach of Hall & Vial (2006) and Bathia et al. (2010), the description length approach of Poskitt & Sengarapillai (2013), pseudo-AIC (Shibata, 1981), scree plot (Cattell, 1966), the eigenvalue ratio criterion (Li et al., 2020), and eigenvector variability plot (Tu et al., 2009). The functional time-series forecasting method is robust to over-estimating $K$, but under-estimating $K$ can result in poor performance. The forecasting method is robust to over-estimating $K$, but under-estimating $K$ can result in inferior accuracy. As in Hyndman et al. (2013), we select $K=6$.
@@
where $\widehat{\beta}_{n+h|n,k}$ denotes the $h$-step-ahead forecast of the principal component scores obtained from a univariate time series forecasting method, such as the autoregressive integrated moving average (ARIMA) model.
@@ 4 Forecasting the Australian sub-national population
To obtain the CWR, that is age 0–1, the ARIMA model is used to model the CWR time series from 1971–2021. Similarly, a one-year ahead forecast for the year 2022 is obtained using automatically selected ARIMA models. Then, the infant age group for males and females can be obtained by (1). By combining the 2022 CWR and CCR forecast and multiplying the population in 2021, we can obtain the population forecast for 2022. Therefore, incorporating the population forecast into the original dataset, the CCR and CWR of 2022 can be calculated accordingly. Incorporating the forecast population into the original population and repeating this process 30 times, we obtain the age-specific population forecast for 2022–2051. By aggregating the sub-national population forecasts, we obtain the population size forecast by sex from 2022 to 2051 in Figure 4. It indicates that the trend for a population for different age groups in the next 30 years varies. For example, the infant age group for females is increasing, but the infant age group for males is decreasing.
@@
Figure 5 depicts the historical cohort change ratios (CCR) for the Australian female and male populations, respectively, from 1972 to 2021. The graphs use a grey gradient to represent the historical data across the age spectrum from 1 to 100+ years, with the final age group capturing all ages above 100. The projected CCR for the next 30 years (2022-2051) is also shown in a vibrant rainbow color scheme. Upon analysing these plots, it can be observed that both genders depict a relatively stable crude birth rate (CCR) of around 1.0 for most age groups, indicating a consistent population size from one generation to the next within these cohorts. The colors of the plots transition smoothly across the spectrum, which indicates a continuum of time. The grey gradient blends into the colors of the rainbow while moving from historical data to future forecasts. These visualizations offer valuable insights into long-term demographic trends and can significantly aid in resource planning and policy formulation in response to changing population dynamics.
@@
Figure 5: Rainbow plots of Australian female and male CCR from 1972 to 2021 (grey gradient color) and their forecasts from 2022 to 2051 (rainbow color) between ages 1 and 100+ in single years of age, with the last age group being 100+.
@@
Figure 6 shows a solid red line representing the forecast female and male population for 2051 (a 30-year forecast). In contrast, the blue dashed line represents the population baseline for 2021. The grey area within the figure represents all the possible scenarios within the 95% pointwise prediction interval for the simulated population paths for 2051. Observing the figure, it is noticeable that both the male and female populations exhibit similar prediction intervals. However, the male population may experience a faster growth rate than the female population.
(a) Female population in Australia
(b) Male population in Australia
@@
Figure 6: Population for 2021 and its forecast for 2051 for each sex, along with 95% pointwise prediction intervals.
@@
Figure 7 displays the forecast total population for males and females. The line depicted in the graph illustrates the current total population for both genders. In contrast, the dashed line within the grey area represents the estimated total population from 2022 to 2051. Furthermore, the grey area represents all possible results within 95% prediction intervals for the total population.
(a) Total female population in Australia
(b) Total male population in Australia
@@
Figure 7: 30-year forecasts of the total population for each sex, along with 95% pointwise prediction intervals.
@@
Figure 8 presents the population pyramids of the forecast population from 2022 to 2051, produced by the functional time series forecast method, with the population in 2021 as a baseline. Figures 7 and 8 indicate that there will be a significant rise in the number of males and females in the next 30 years. The male population is expected to grow more rapidly than the female population during this period. Population growth is a crucial factor that the Federal Government needs to track to ensure that the community is prepared and equipped with the necessary resources. The forecast is vital to make informed decisions and take preventive measures to meet the future needs of the growing population.
@@
Figure 9 presents a comparative graph analysis of the total population trends by gender for various Australian states and territories from 1972 to 2051. Figures 9a and 9b delineate the population trajectories for females and males, respectively. Each line within the graph corresponds to a different state, such as ACT, NSW, VIC and so on.
@@
No abrupt discontinuities are evident within the presented data range, suggesting a steady population increase rather than sudden demographic shifts. The implications of these trends are significant for policy planning, particularly in determining sustainable pension ages. The gradual population increase, especially within the working-age cohort, will directly affect the pension system, necessitating adjustments to pension ages and related social security frameworks to ensure long-term sustainability.
Figure 10 exhibits a graphical display representing each Australian state’s total population across a temporal span from 1972 to 2051. The graph is not segmented by gender, it provides an aggregate overview of demographic trends within each region. Similarly to Figure 9, each line on the graph represents a different state or territory, with the same labels from Figure 9 serving as identifiers. The lines depict a clear upward progression in population numbers, indicating overall growth within these regions. The slope of each line varies, suggesting that the rate of population increase is not uniform across the states and territories. Some states, such as NSW and VIC, are typically populous. They display a steeper incline, indicative of a more rapid population rise.

### v2

@@ Abstract
The age pension aims to assist eligible elderly Australians who meet specific age and residency criteria in maintaining basic living standards. In designing efficient pension systems, government policymakers seek to satisfy the expectations of the overall aging population in Australia. However, the population’s unique demographic characteristics at the state and territory level are often overlooked due to the lack of available data. We use the Hamilton-Perry model, which requires minimum input, to model and forecast the evolution of age-specific populations at the state level. We also integrate the obtained sub-national demographic information to determine sustainable pension ages up to 2051. We also investigate pension welfare distribution in all states and territories to identify disadvantaged residents under the current pension system. Using the sub-national mortality data for Australia from 1971 to 2021 obtained from AHMD (2023), we implement the Hamilton-Perry model with the help of functional time series forecasting techniques. With forecasts of age-specific population sizes for each state and territory, we compute the old age dependency ratio to determine the nationwide sustainable pension age.
@@ 1 Introduction
We propose modeling and projecting state and territory populations using the Hamilton-Perry (HP) model (Hamilton & Perry, 1962) and use the forecasts to determine the sustainable pension age. The HP model overcomes the restriction of data availability at the sub-national level since it requires significantly fewer inputs and computational time to implement than the traditional cohort-component method (see, e.g., Shang et al., 2016) or multiregional population projection (see, e.g., Rogers, 1975). Specifically, the standard cohort-component approach for population projection requires high-quality data about mortality, fertility, and immigration of a population and assumptions. In contrast, the HP model only requires population size by age and sex from the two most recent censuses (Smith et al., 2001, Chapter 7). This has made the HP model popular over the last decade in small area population forecasts, with notable applications including an initial population projection for Clark County, Nevada (Swanson et al., 2010) and a subsequent nationwide census tract evaluation across all sub-county areas in the United States (Baker et al., 2021), to name only a few. Most recently, a modified HP model was considered to forecast populations for small areas of Australia at the statistical area Level 2 scale (defined as in Australian Bureau of Statistics, 2011). This modified HP model with smoothed age profiles produces the most accurate forecasts (Wilson & Grossman, 2022).
This paper contributes a new variant of the HP model by extending the idea of a smoothing cohort change ratio (CCR, Wilson & Grossman, 2022) in combination with functional time series forecasting techniques (see Ramsay & Silverman, 2005; Ferraty & Vieu, 2006, for an introduction to functional data analysis). Functional time series methods have been widely used in demographic data modeling and forecasting (see, e.g., Cairns et al., 2011; D’Amato et al., 2011; Carfora et al., 2017). Using this framework, the CCR of sex $s$ increasing from age $x$ to $x+1$ within the time interval $[t-1,t]$, $\text{CCR}^{s}_{x\rightarrow x+1}(t-1\rightarrow t)$, can be regarded as a stochastic process defined on a compact set $x\in\mathcal{I}$ and indexed by a time label $t\in\mathbf{Z}$. This functional time series setting is best known for retaining the correlation of demographic observations at different ages, which enables extracting the temporal dependence of the underlying stochastic process for forecasting purposes. These advantages allow functional time series methods to produce more accurate forecasts than the commonly used parametric models (see, e.g., Hyndman & Ullah, 2007). With minimal data input, we combine functional time series forecasting techniques with the HP model to produce accurate sub-national population forecasts. By aggregating the sub-national population forecasts to the national level, we determine a sustainable pension age based on the current old-age dependency ratio.
@@
The remainder of the paper is organized as follows. Section 2 describes the age- and sex-specific population data at Australia’s national and sub-national levels. Section 3 provides details of the population forecasting method. The empirical data analysis results are presented in Section 4. Section 5 proposes an approach to determine the optimal pension age based on the forecasted population and its old-age dependency ratio. With the help of the HP model, we study the changes in population at the sub-national level. From the historical observed data, we generate out-of-sample population forecasts from 2022 to 2051 and determine a sustainable pension age for Australia. With the sub-national population projections and forecasts of pension payout rates, we compare lifetime age pension income for residents in Australia’s six states and two territories in Section 6. Finally, Section 7 concludes the paper and provides some directions for future research.
@@ 2.2 Sub-national age- and sex-specific data
While national-level data provides an overall picture of a country’s demographic trends, economic indicators, and other key statistics, sub-national details are crucial for understanding regional variations and making more targeted policy decisions. Sub-national data allow for granular analysis of regional disparities, urban-rural differences, resource allocation, and infrastructure planning.
@@ 3.1 Hamilton-Perry model
A recent study by Wilson & Rees (2021) outlined the standard process for creating a national population forecast using a cohort-component model. This model requires high-quality data on fertility, mortality, and internal and international migration. The quality of these datasets needs to be carefully examined to ensure accurate population forecasts. The cohort-component model is advantageous as it effectively represents underlying population processes and provides age-disaggregated projections for various purposes such as government planning, policy-making, budgeting, and service delivery. However, it can also be highly inaccurate if poor assumptions are made about fertility, mortality, and migration. Additionally, the cohort-component model requires significant input data, and assumptions about future fertility, mortality, and migration must be prepared separately.
Compared to the cohort-component model, Hamilton & Perry’s (1962) method overcomes these drawbacks as it requires much smaller data requirements for projection. Instead of modeling mortality, fertility, and migration, the HP model requires data from the two most recent censuses (Smith et al., 2001, pp. 153–158). The ratio between two population censuses is known as the CCR. It measures the combined net effect of mortality and migration experienced by the cohort over time. It is estimated from population sizes by age and sex for the jump-off year and (commonly) one year earlier. For example, the CCR for the cohort of sex $s$ at age interval $x$ at time $t-1$ and age $x+1$ at time $t$ is defined as
@@
The HP model is commonly used for predicting the population of small areas (e.g., Wilson et al., 2022). With the sub-national population forecasts for eight regions, we study the dynamic changes in regional population counts. By aggregating the population forecasts at the national level, we provide Australian Government policymakers with a future sustainable retirement age.
@@ 3.2 Functional time series forecasting
[…]
@@
The $K(u,v)$ is an analogue of the variance-covariance matrix commonly used in multivariate data analysis. The variation in population functions over the entire domain of the age variable $u$ is captured by the variance function (Ramsay & Silverman, 2005, Chapter 8).
@@
There are many ways of determining the optimal number of $K$, such as the bootstrap approach of Hall & Vial (2006) and Bathia et al. (2010), the description length approach of Poskitt & Sengarapillai (2013), pseudo-AIC (Shibata, 1981), scree plot (Cattell, 1966), the eigenvalue ratio criterion (Li et al., 2020), and eigenvector variability plot (Tu et al., 2009). The functional time-series forecasting method is robust in overestimating $K$, but underestimating $K$ can result in poor performance. The forecasting method is robust in overestimating $K$, but underestimating $K$ can result in inferior accuracy. As in Hyndman et al. (2013), we select $K=6$.
@@
where $\widehat{\beta}_{n+h|n,k}$ denotes the $h$-step-ahead forecast of the principal component scores obtained from a univariate time series forecasting method, such as the autoregressive integrated moving average (ARIMA) model. In our application, the described process forecasts the mortality function $\mathcal{X}_{t}(u)$ for the entire age range $u\in[0,100+]$. Consequently, the interdependence among different ages is preserved during the forecasting.
@@ 3.3 Construction of prediction intervals via nonparametric bootstrapping
3.4 Multivariate functional time series method
Inspired by joint modeling, a multivariate functional time series method is proposed to model and forecast multiple series simultaneously (Shang & Kearney, 2022). The multivariate functional time series are stacked in a vector with $\bm{\mathcal{X}}_{t}(u)=[\mathcal{X}_{t}^{\text{F}}(u),\mathcal{X}_{t}^{\text{M}}(u)]$. For $u,v\in\mathcal{I}$, the theoretical cross-covariance function can be defined with elements
$c_{\text{F, M}}(u,v):$ $=\text{Cov}[\mathcal{X}^{\text{F}}(u),\mathcal{X}^{\text{M}}(u)]$
$=\text{E}\left\{[\mathcal{X}^{\text{F}}(u)-\mu^{\text{F}}(u)][\mathcal{X}^{\text{M}}(u)-\mu^{\text{M}}(u)]\right\}$
where $\mu^{\text{F}}(u)$ and $\mu^{\text{M}}(u)$ denote the mean functions for the female and male series, respectively. Via Mercer’s lemma, there exists an orthonormal sequence $(\phi_{k})$ and a non-increasing sequence $\lambda_{k}$ of positive numbers, such that
$c_{\text{F, M}}(u,v)=\sum^{\infty}_{k=1}\lambda_{k}\phi_{k}^{\text{F}}(u)\phi_{k}^{\text{M}}(v).$
By Karhunen-Loève expansion, a realization of the stochastic process $\bm{\mathcal{X}}(u)$ can be expressed as
$\mathcal{X}_{t}^{\text{F}}(u)=\overline{\mathcal{X}}^{\text{F}}(u)+\sum^{K}_{k=1}\widehat{\beta}_{t,k}^{\text{F}}\widehat{\phi}_{k}^{\text{F}}(u),$
where $K$ denotes the number of retained functional principal components which is the same across multiple series, $\widehat{\phi}_{k}^{\text{F}}(u)$ represents the $k$th estimated functional principal component corresponding to the female population, and $\widehat{\beta}_{t,k}^{\text{F}}=\langle\mathcal{X}_{t}^{\text{F}}(u)-\overline{\mathcal{X}}^{\text{F}}(u),\phi_{k}^{\text{F}}(u)\rangle$ represents $k$th principal component score for the female population in year $t$. A similar decomposition can also be implemented for the male series.
Conditioning on the past curves $\bm{\mathcal{X}}(u)=[\bm{\mathcal{X}}_{1}(u),\dots,\bm{\mathcal{X}}_{n}(u)]$ for both series and the estimated functional principal components $\bm{\Phi}(u)=\big{[}\widehat{\bm{\phi}}^{\text{F}}(u),\widehat{\bm{\phi}}^{\text{M}}(u)\big{]}$, the $h$-step-ahead point forecast of $\bm{\mathcal{X}}_{n+h}(u)$ can be expressed as
$\widehat{\bm{\mathcal{X}}}_{n+h|n}(u)$ $=\text{E}[\bm{\mathcal{X}}_{n+h}(u)|\bm{\mathcal{X}}(u),\bm{\Phi}(u)]$
$=\overline{\bm{\mathcal{X}}}(u)+\bm{\Phi}(u)\widehat{\bm{\beta}}_{n+h|n}^{\top},$
where $\overline{\bm{\mathcal{X}}}(u)=\big{[}\overline{\mathcal{X}}^{\text{F}}(u),\overline{\mathcal{X}}^{\text{M}}(u)\big{]}^{\top}$ denotes the mean function aggregated over two populations, and let $\widehat{\bm{\beta}}_{n+h|n}^{\top}=\big{[}\widehat{\bm{\beta}}_{n+h|n}^{\text{F}},\widehat{\bm{\beta}}_{n+h|n}^{\text{M}}\big{]}$ denote the $h$-step-ahead forecasts of the principal component scores obtained from a univariate time-series forecasting method.
@@ 4 Forecasting the Australian sub-national population
To obtain the CWR, that is, age 0–1, the ARIMA model is used to model the CWR time series from 1971–2021. Similarly, a one-year ahead forecast for the year 2022 is obtained using automatically selected ARIMA models. Then, the infant age group for males and females can be obtained by (1). By combining the 2022 CWR and CCR forecast and multiplying the population in 2021, we can obtain the population forecast for 2022. Therefore, incorporating the population forecast into the original dataset, the CCR and CWR of 2022 can be calculated accordingly. Incorporating the forecast population into the original population and repeating this process 30 times, we obtain the age-specific population forecast for 2022–2051. By aggregating the sub-national population forecasts, we obtain the population size forecast by sex from 2022 to 2051 in Figure 4. It indicates that the trend for a population for different age groups in the next 30 years varies. For example, the infant age group for females is increasing, but the infant age group for males is decreasing.
@@
The functional time series approach provides a robust population forecast by treating male and female populations as independent variables. However, this method does not account for the correlation between the two sexes, potentially overlooking important dynamics. To address this limitation, we introduce the “Multivariate Functional Time Series” (MFTS) method.
The MFTS method enhances forecasting accuracy by capturing the correlation between male and female populations. It does this by combining the Coherent Compositional Reconciliation (CCR) for both sexes, treating them as dependent variables. By incorporating the interdependence between male and female populations, the MFTS method offers a more comprehensive and nuanced forecast. This approach ensures that changes in one sex’s population are appropriately reflected in the other, providing a holistic view of population dynamics. As a result, the MFTS method complements univariate functional time series forecasting, delivering a more detailed and accurate prediction of future population trends.
Figure 5 depicts the historical cohort change ratios (CCR) for the Australian female and male populations, respectively, from 1972 to 2021. The graphs use a grey gradient to represent the historical data across the age spectrum from 1 to 100+ years, with the final age group capturing all ages above 100. The projected CCR for the next 30 years (2022-2051) is also shown in a vibrant rainbow color scheme. Upon analyzing these plots, it can be observed that both genders depict a relatively stable crude birth rate (CCR) of around 1.0 for most age groups, indicating a consistent population size from one generation to the next within these cohorts. The colors of the plots transition smoothly across the spectrum, which indicates a continuum of time. The grey gradient blends into the colors of the rainbow while moving from historical data to future forecasts. These visualizations offer valuable insights into long-term demographic trends and can significantly aid resource planning and policy formulation in response to changing population dynamics.
@@
Figure 5: Rainbow plots of Australian female and male CCR from 1972 to 2021 (grey gradient color) and their forecasts from 2022 to 2051 (rainbow color) between ages one and 100+ in single years of age, with the last age group being 100+.
@@
Figure 6 shows a solid red line representing the forecast female and male population for 2051 (a 30-year forecast). In contrast, the blue dashed line represents the population baseline for 2021. The grey area within the figure represents all the possible scenarios within the 95% pointwise prediction interval for the simulated population paths for 2051. Observing the figure, it is noticeable that both the male and female populations exhibit similar prediction intervals. However, the male population may experience a faster growth rate than the female population. Notably, the prediction intervals for both males and females are more refined in Figures 6 (c) and (d), demonstrating the MFTS method’s capability to account for the correlation between sexes. However, the differences between the univariate and the multivariate functional time series methods are small. Although the MFTS method provides a distinct forecast, the overall influence on future population trends is similar. This suggests that while the MFTS method offers a different perspective by considering the interdependence between male and female populations, the general trends and impacts on future population projections remain consistent across both methods.
(a) Univariate functional time series method
(b) Univariate functional time series method
(c) Multivariate functional time series method
(d) Multivariate functional time series method
@@
Figure 6: Population for 2021 and its forecast for 2051 for each sex, along with 95% prediction intervals.
@@
Figure 7 displays the forecast total population for males and females. The line depicted in the graph illustrates the current total population for both genders. In contrast, the dashed line within the grey area represents the estimated total population from 2022 to 2051. Furthermore, the grey area represents all possible results within 95% prediction intervals for the total population. Figures 7 (c) and (d) present the forecasts using the MFTS method. The blue line represents actual population data up to 2021, while the red line shows the forecasted population from 2022 to 2051. Notably, the point forecasts using the MFTS method are slightly lower than those obtained using the functional time series method. Additionally, the grey-shaded areas representing the 95% prediction intervals are slightly wider in the MFTS forecasts.
(a) Univariate functional time series method
(b) Univariate functional time series method
(c) Multivariate functional time series method
(d) Multivariate functional time series method
@@
Figure 7: 30-year forecasts of the total population for each sex, along with 95% prediction intervals.
This suggests that while the MFTS method accounts for more uncertainty, it provides a comprehensive view by including the correlation between male and female populations, potentially offering a more nuanced understanding of future population trends. Overall, Figure 7 highlights the differences and similarities between the two forecasting methods, demonstrating that although the point forecasts differ slightly, the general trends and prediction intervals remain broadly consistent. This comparison underscores the value of considering multiple methods to capture a range of potential future scenarios.
@@
Figure 8 presents the population pyramids of the forecast population from 2022 to 2051, produced by the functional time series forecast method, with the population in 2021 as a baseline. Figures 7 and 8 indicate that there will be a significant rise in the number of males and females in the next 30 years. The male population is expected to grow more rapidly than the female population during this period. Population growth is a crucial factor that the Federal Government needs to track to ensure that the community is prepared and equipped with the necessary resources. Forecasting is vital to making informed decisions and taking preventive measures to meet the future needs of the growing population.
@@
Figure 9 presents a comparative graph analysis of the total population trends by gender for various Australian states and territories from 1972 to 2051. Figures 9a and 9b delineate the population trajectories for females and males, respectively. Each line within the graph corresponds to a different state, such as ACT, NSW, VIC, and so on.
@@
No abrupt discontinuities are evident within the presented data range, suggesting a steady population increase rather than sudden demographic shifts. These trends have significant implications for policy planning, particularly in determining sustainable pension ages. The gradual population increase, especially within the working-age cohort, will directly affect the pension system, necessitating adjustments to pension ages and related social security frameworks to ensure long-term sustainability.
Figure 10 exhibits a graphical display representing each Australian state’s total population across a temporal span from 1972 to 2051. The graph is not segmented by gender, and it provides an aggregate overview of demographic trends within each region. Similarly to Figure 9, each line on the graph represents a different state or territory, with the same labels from Figure 9 serving as identifiers. The lines depict a clear upward progression in population numbers, indicating overall growth within these regions. The slope of each line varies, suggesting that the rate of population increase is not uniform across the states and territories. Some states, such as NSW and VIC, are typically populous. They display a steeper incline, indicative of a more rapid population rise.
@@
In Figure 11, we display the population forecasts for Australia and the combination of forecasts from its subnational level. Two different population forecasting models were fitted to the population data from 1971 to 2021 to produce forecasts from 2022 to 2051. The forecasting model based on the Australian national data can smooth out regional variations and anomalies, providing a broader but less detailed perspective. On the other hand, the forecasting model using subnational population data considers regional specification, capturing local trends and demographic changes, thus offering more accurate and detailed predictions.
Figure 11: Comparison of national vs. sub-national population forecasts for Australia (2022-2051).
From Figure 11, it reveals that the two forecast lines may diverge, indicating differences in prediction when using aggregated national data versus detailed subnational data. The subnational forecasts in red show high variability, emphasizing the importance of considering regional differences in making population forecasts. The divergence of the lines over time can indicate changing regional dynamics that might not be apparent when studying at national data alone.